OpenStudy (anonymous):

Prove:

OpenStudy (anonymous):

\[\frac{ \sin^4x-\cos^4x }{ 1-\cot^4x }=\sin^4x\]

OpenStudy (anonymous):

\[\sin^4 x-\cos^4 x=(\sin^2 x +\cos^2 x)(\sin^2 x - \cos^2 x) = \sin^2 x-\cos^2 x\]

OpenStudy (anonymous):

\[1-\cot^4 x = (1-\cot^2 x)(1+\cot^2 x) = \csc^2 x (1-\cot^2 x)\]

OpenStudy (anonymous):

my internet is working slow .....I cant see the equations u have written properly

OpenStudy (anonymous):

can u draw?

OpenStudy (anonymous):

Use the identity a^2-b^2 = (a-b)(a+b)

OpenStudy (anonymous):

\[\frac{\sin^4x-\cos^4x}{1-\cot^4x}=\frac{\sin^4x-\cos^4x}{1-\frac{\cos^4x}{\sin^4x}}\\ =\frac{\sin^4x-\cos^4x}{\frac{\sin^4x-\cos^4x}{\sin^4x}}=\sin^4x \]

OpenStudy (anonymous):

It is rather pointless if one just give the answer right away.

OpenStudy (anonymous):

I cannot see any of the equations listed above ^^

OpenStudy (anonymous):

plz can someone draw....

OpenStudy (anonymous):

the answer is right there.. just notice that the answer only has the sin^4x term. so, what is the factor on the other side that is causing the trouble? the cot.. so, convert it and viola

OpenStudy (anonymous):

Nope, you see, we have \[\frac{\sin^2 x}{\csc^2 x}\]

OpenStudy (anonymous):

:(

OpenStudy (zehanz):

Here is a screenshot of the formulas typed in earlier: (hope you can read them)

OpenStudy (anonymous):

if you factorize it, you decrease the order of the system. you final answer is still in sin^4 so, factorization will cause trouble.

OpenStudy (anonymous):

@ZeHanz thanks it was kindof blind situation for me :D

OpenStudy (zehanz):

@koli123able: It took a while for us to realize that :)

OpenStudy (anonymous):

how did (sin^2x+cos^2x)(sin^2x-cos^2x) = sin^2x-cos^2x ??

OpenStudy (anonymous):

@agostino ?

OpenStudy (anonymous):

@ZeHanz the screenshot says --> (sin^2x+cos^2x)(sin^2x-cos^2x) = sin^2x-cos^2x how did this happen??

OpenStudy (zehanz):

It is because sin²x+cos²x=1 for all x.

OpenStudy (anonymous):

what to do after sin^2x-cos^x/cosec^x(1-cot^2x)

OpenStudy (zehanz):

@koli123able: although @electrokid received some critique about his giving away the answer, it is in fact the right one: All this factoring and converting is a little over the top. Why? Because what you have to prove is that the left hand side is equal to the rhs. Everything there is about 4th powers. So there is no need to factor, and get lower than 4th powers. Have a good look at his calculation and you will see what it is about. To be able to see it, I'll include another screenshot :D

OpenStudy (anonymous):

@koli123able and @agostino you are welcome @ZeHanz Thanks

OpenStudy (zehanz):

YW!